OpenStudy (anonymous):

the sum of a number and its reciprocal is 13. find the number?

OpenStudy (anonymous):

so its x+ 1/x= 13

OpenStudy (anonymous):

but what do i do from there?

OpenStudy (anonymous):

add and get \[\frac{x^2+1}{x}=13\]

OpenStudy (anonymous):

then write \[x^2+1=13x\] so \[x^2-13x+1=0\] and solve for \(x\)

OpenStudy (anonymous):

then use the quadratic formula you good from there?

OpenStudy (anonymous):

yeah, i used the quadratic formula and got 12.92, -.07 so i figured i was wrong somehow

OpenStudy (anonymous):

i would go with \[\frac{13+\sqrt{165}}{2}\]

OpenStudy (anonymous):

haha okay, thanks again!!!